A baseball team has home games on Friday and Saturday. The two games together earn $3853.50 for the team. Friday's game generates $486.50 less than Saturday's game. How much money was taken in at each game?
Answer by Lstark01(4)

Let F = Friday game
Let S = Saturday game
Your 2 equations are going to be:
F + S = 3853.50
F = S - 456.50
Since F = S - 456.50 already has one equation solved, use the substitution method to solve.
(S - 456.50) + S = 3853.50 (combine like terms)
2S - 456.50 = 3853.50
Add 456.50 to both sides
2S = 4310.00
Divide 2 on both sides to solve for S
S = 2155.00
Use S = 2155.00 to solve for F.
F + 2155.00 = 3853.50
Subtract 2155.00 from both sides
F = 1698.50
Depending on the teacher, you might need to answer in a complete sentence. So you would say:
Fridays game generated $1698.50 and Saturdays game generated $2155.00.